# Critical scaling profile for trees and connected subgraphs on the complete graph

### v1

@@ 1.1 Results
The enumeration of trees and connected graphs has a long history. We are motivated by problems arising in the critical behaviour of branched polymers in equilibrium statistical mechanics, to consider certain generating functions for the number of trees and connected subgraphs in the complete graph $\mathbb{K}_{V}$ on $V$ labelled vertices. The vertices are labelled as $\mathbb{V}=\{0,\dots,V-1\}$ and the edge set is $\mathbb{E}=\{\{x,y\}:x,y\in\mathbb{V}\}$. Our interest is in the asymptotic behaviour as $V\to\infty$.
@@
where the first sum is over all labelled trees $T$ in $\mathbb{K}_{V}$ containing the vertex 00, the second sum is over all labelled connected subgraphs containing 00, and $|T|$ and $|A|$ denote the number of edges in $T$ and $A$. The division of $p$ by $eV$ is a normalisation to make $p=1$ correspond to a critical value. We also study the two-point functions
@@ 1.2.1 Trees
Similarly,
@@ 1.2.2 Connected subgraphs
Note that terms in the above sum are zero unless $\ell\leq\binom{n}{2}-(n-1)$, and that the tree term ($\ell=0$) is absent. We extract the tree contributions and define remainder terms by
@@ 1.3 Motivation
$g(z)=\sum_{m=0}^{\infty}c_{m}(0)z^{m},\qquad G_{z}(x)=\sum_{m=0}^{\infty}c_{m}(x)z^{m},\qquad\chi(z)=\sum_{x\in\mathbb{Z}^{d}}G_{z}(x).$ (1.24)
@@
In (1.25), the torus point $x$ is identified with its representative in $\mathbb{Z}^{d}\cap(-\frac{r}{2},\frac{r}{2}]^{d}$ in the evaluation of $G_{z_{c}}(x)$. For $d>8$, $G_{z_{c}}(x)$ decays as $|x|^{-(d-2)}$ [10, 11], and the constant term of order $V^{-3/4}=r^{-3d/4}$ dominates the Gaussian decay over most of the torus. This is the “plateau” phenomenon. On the complete graph, only the plateau term occurs for the two-point function at distinct points, as in (1.6). For $d=d_{\rm c}=8$, the conjecture is modified to include logarithmic corrections to the window scale $V^{-1/2}$, the plateau scale $V^{-3/4}$, and the susceptibility scale $V^{1/4}$, but with the identical profile $I_{0}$.
Self-avoiding walk: Self-avoiding walk on the complete graph $\mathbb{K}_{V}$ is exactly solvable [27]. Let $c_{V,n}(0,1)=\prod_{j=1}^{n}(V-j)$ denote the number of $n$-step self-avoiding walks from 00 to $1$ on $\mathbb{K}_{V}$. Let $S_{V,01}(p)=\sum_{n=1}^{\infty}c_{V,n}(0,1)(p/V)^{n}$ and let $\chi_{V}^{\rm SAW}(p)=1+(V-1)S_{V,01}(p)$. It is proved in [27] (see also [24, Appendix B]) that, as $V\to\infty$,
@@
in distribution [1]. The case $k=2$ case corresponds to $\chi_{V}^{\rm{perc}}$ and identifies $f_{\rm perc}(s)$.
@@ 2 Proof for trees
$\sum_{n=b\sqrt{V}}^{V}\frac{1}{n^{\gamma}}e^{-\kappa n^{2}/V}e^{\lambda n/\sqrt{V}}\leq C_{\kappa,\lambda}b^{-\gamma}V^{(1-\gamma)/2}.$ (2.1)
@@
$\frac{1}{b^{\gamma}V^{\gamma/2}}\sum_{n=b\sqrt{V}}^{\infty}e^{-\kappa n^{2}/V}e^{\lambda n/\sqrt{V}}.$ (2.2)
@@
$\chi_{V}^{(1)}\lesssim\sum_{n=1}^{a\sqrt{V}}1\frac{1}{\sqrt{n}}e^{|s|n/\sqrt{V}}\leq e^{a\lvert s\rvert}\sum_{n=1}^{a\sqrt{V}}\frac{1}{\sqrt{n}}\lesssim e^{a\lvert s\rvert}(1+a^{1/2}V^{1/4}),$ (2.9)
as claimed. For $\chi_{V}^{(3)}$, we also need a bound on the product over $j$. Using $\log(1+x)\leq x$, we have
$\prod_{j=1}^{n-1}\big{(}1-\frac{j}{V}\big{)}=\exp\Bigl{\{}\sum_{j=1}^{n-1}\log(1-\frac{j}{V})\Bigr{\}}\leq\exp\Bigl{\{}-\frac{1}{V}\sum_{j=1}^{n-1}j\Bigr{\}}=\exp\Bigl{\{}-\frac{n(n-1)}{2V}\Bigr{\}}.$ (2.10)
@@
$\chi_{V}^{(3)}\lesssim\sum_{n=b\sqrt{V}}^{V}e^{-n^{2}/2V}e^{n/2V}\frac{1}{\sqrt{n}}e^{|s|n/\sqrt{V}}\lesssim e^{1/2}b^{-1/2}V^{1/4}.$ (2.11)
@@
which follow from Taylor expansion of the logarithm. Since $n\in[a\sqrt{V},b\sqrt{V}]$, the above, together with the fact that $b_{n}=\frac{e}{\sqrt{2\pi n}}[1+O(1/n)]$ by Stirling’s formula, give
$\chi_{V}^{(2)}$ $=\sum_{n=a\sqrt{V}}^{b\sqrt{V}}e^{-n^{2}/2V}\frac{e}{\sqrt{2\pi n}}e^{sn/\sqrt{V}}\Big{[}1+O\Big{(}\frac{1}{\sqrt{V}}\Big{)}\Big{]}.$ (2.14)
@@
$G_{V}^{(2)}(p_{V})\leq G_{V}^{(2)}(1+sV^{-1/2})\leq e^{bs}\sum_{n=N}^{b\sqrt{V}}\frac{1}{n^{3/2}}\lesssim\frac{e^{bs}}{\sqrt{N}}.$ (2.17)
@@ 3.1 Bound on $C(n,n+k)$
We use the asymptotic formula (3.5), and use that $w_{k}=1+O(1/k)$ is bounded. The error term in (3.5) is bounded by a constant since $k\leq N-n\leq n^{2}$. The factor $e^{a(x)}$ is also bounded by a constant, by Lemma 3.1(i). We therefore only need to estimate $e^{n\varphi(x)}$. Since $0\leq\frac{k}{n}\leq\frac{t^{2}}{3e}$ and $x=1+\frac{k}{n}$, by (3.2) we have $y(x)\leq\sqrt{3(x-1)}=\sqrt{3k/n}\leq t/\sqrt{e}$, so Lemma 3.1(ii) gives
@@ 3.2 Bound on the surplus generating function
$A(n,z)$ $=\frac{1}{n^{n-2}}\sum_{k=0}^{n}C(n,n+k)z^{k+1},\qquad B(n,z)=\frac{1}{n^{n-2}}\sum_{k=\frac{1}{2}n}^{\infty}C(n,n+k)z^{k+1},$ (3.15)
@@ 3.3 Proof for connected subgraphs
$\Delta_{V}^{(1)}$ $\leq C_{b}V^{-1/4}\sum_{n=3}^{b\sqrt{V}}\binom{V-1}{n-1}\Big{(}\frac{p}{eV}\Big{)}^{n-1}n^{n-1}\leq C_{b}V^{-1/4}\chi_{V}^{t}(p).$ (3.31)
@@
$\Delta_{V}^{(2)}(1+sV^{-1/2})\lesssim\sum_{n=b\sqrt{V}}^{V}\binom{V-1}{n-1}\bigg{(}\frac{1+sV^{-1/2}}{eV}\bigg{)}^{n-1}n^{n-1}\exp\Bigl{\{}\frac{1}{5}\frac{n^{3}}{V^{2}}\Bigr{\}}.$ (3.34)
@@
$\Delta_{V}^{(2)}(1+sV^{-1/2})\lesssim\sum_{n=b\sqrt{V}}^{V}e^{-(\frac{1}{2}-\frac{1}{5})n^{2}/V}\frac{1}{\sqrt{n}}e^{|s|n/\sqrt{V}}\leq C_{\lvert s\rvert}b^{-1/2}V^{1/4}.$ (3.35)
@@
$\Delta_{V,0}^{(2)}(1+sV^{-1/2})\lesssim\sum_{n=b\sqrt{V}}^{V}e^{-(\frac{1}{2}-\frac{1}{5})n^{2}/V}\frac{1}{n^{3/2}}e^{|s|n/\sqrt{V}}\leq C_{\lvert s\rvert}b^{-3/2}V^{-1/4}.$ (3.38)

### v2

@@ 1.1 Results
The enumeration of trees and connected graphs has a long history. We are motivated by problems arising in the critical behaviour of branched polymers in equilibrium statistical mechanics, to consider certain generating functions for the number of trees and connected subgraphs in the complete graph $\mathbb{K}_{V}$ on $V$ labelled vertices. The vertices are labelled as $\mathbb{V}=\{0,\dots,V-1\}$ and the edge set is $\mathbb{E}=\{\{x,y\}:x,y\in\mathbb{V},\ x\neq y\}$. Our interest is in the asymptotic behaviour as $V\to\infty$.
@@
where the first sum is over all labelled trees $T$ in $\mathbb{K}_{V}$ containing the vertex 00, the second sum is over all labelled connected subgraphs $A$ containing 00, and $|T|$ and $|A|$ denote the number of edges in $T$ and $A$. The division of $p$ by $eV$ is a normalisation to make $p=1$ correspond to a critical value. We also study the two-point functions
@@ 1.2.1 Trees
Similarly, by counting the number of ways to choose $n-2$ vertices other than 00 and $1$, we have
@@ 1.2.2 Connected subgraphs
Note that terms in the above sum are zero unless $\ell\leq\binom{n}{2}-(n-1)$, and that the tree term ($\ell=0$) is absent.
We decompose the sums defining $G_{V,0}^{a}$ and $G_{V,01}^{a}$ according to the number $n$ of vertices in the connected subgraph, and we further distinguish whether or not the subgraph contains surplus edges. This leads to the decomposition
@@ 1.3 Motivation
$g(z)=\sum_{m=0}^{\infty}c_{m}(0)z^{m},\quad G_{z}(x)=\sum_{m=0}^{\infty}c_{m}(x)z^{m},\quad\chi(z)=\sum_{x\in\mathbb{Z}^{d}}G_{z}(x).$ (1.24)
@@
In (1.25), the torus point $x$ is identified with its representative in $\mathbb{Z}^{d}\cap(-\frac{r}{2},\frac{r}{2}]^{d}$ in the evaluation of $G_{z_{c}}(x)$. For $d>8$, $G_{z_{c}}(x)$ decays as $|x|^{-(d-2)}$ [10, 11], and the constant term of order $V^{-3/4}=r^{-3d/4}$ dominates the Gaussian decay over most of the torus. This is the “plateau” phenomenon. On the complete graph, the decaying term $\lvert x\rvert^{-(d-2)}$ is absent, and only the constant term occurs for $G_{V,01}$, as in (1.6). For $d=d_{\rm c}=8$, the conjecture is modified to include logarithmic corrections to the window scale $V^{-1/2}$, the plateau scale $V^{-3/4}$, and the susceptibility scale $V^{1/4}$, but with the identical profile $I_{0}$.
Self-avoiding walk: Self-avoiding walk on the complete graph $\mathbb{K}_{V}$ is exactly solvable [27]. For $11\leq n\leq V-$ 1, let $c_{V,n}(0,1)=\prod_{j=2}^{n}(V-j)$ denote the number of $n$-step self-avoiding walks from 00 to $1$ on $\mathbb{K}_{V}$. Let $S_{V,01}(p)=\sum_{n=1}^{V-1}c_{V,n}(0,1)(p/V)^{n}$ and let $\chi_{V}^{\rm SAW}(p)=1+(V-1)S_{V,01}(p)$. It is proved in [27] (see also [24, Appendix B]) that, as $V\to\infty$,
@@
in distribution [1]. The $k=2$ case corresponds to $\chi_{V}^{\rm{perc}}$ and identifies $f_{\rm perc}(s)$.
@@ 2 Proof for trees
$\sum_{n=\lceil b\sqrt{V}\rceil}^{V}\frac{1}{n^{\gamma}}e^{-\kappa n^{2}/V}e^{\lambda n/\sqrt{V}}\leq C_{\kappa,\lambda}b^{-\gamma}V^{(1-\gamma)/2}.$ (2.1)
@@
$\frac{1}{b^{\gamma}V^{\gamma/2}}\sum_{n=\lceil b\sqrt{V}\rceil}^{\infty}e^{-\kappa n^{2}/V}e^{\lambda n/\sqrt{V}}.$ (2.2)
@@
$\chi_{V}^{(1)}\lesssim\sum_{n=1}^{\lfloor a\sqrt{V}\rfloor}1\frac{1}{\sqrt{n}}e^{|s|n/\sqrt{V}}\leq e^{a\lvert s\rvert}\sum_{n=1}^{\lfloor a\sqrt{V}\rfloor}\frac{1}{\sqrt{n}}\lesssim e^{a\lvert s\rvert}(1+a^{1/2}V^{1/4}),$ (2.9)
as claimed. For $\chi_{V}^{(3)}$, we also need a bound on the product over $j$. Using $1-x\leq e^{-x}$, we have
$\prod_{j=1}^{n-1}\big{(}1-\frac{j}{V}\big{)}\leq\exp\Bigl{\{}-\frac{1}{V}\sum_{j=1}^{n-1}j\Bigr{\}}=\exp\Bigl{\{}-\frac{n(n-1)}{2V}\Bigr{\}}.$ (2.10)
@@
$\chi_{V}^{(3)}\lesssim\sum_{n=\lceil b\sqrt{V}\rceil}^{V}e^{-n^{2}/2V}e^{n/2V}\frac{1}{\sqrt{n}}e^{|s|n/\sqrt{V}}\lesssim e^{1/2}b^{-1/2}V^{1/4}.$ (2.11)
@@
which follow from Taylor expansion of the logarithm (the constants here depend on $s$). Since $n\in[a\sqrt{V},b\sqrt{V}]$, the above, together with the fact that $b_{n}=\frac{e}{\sqrt{2\pi n}}[1+O(1/n)]$ by Stirling’s formula, give
$\chi_{V}^{(2)}$ $=\sum_{n=\lceil a\sqrt{V}\rceil}^{\lfloor b\sqrt{V}\rfloor}e^{-n^{2}/2V}\frac{e}{\sqrt{2\pi n}}e^{sn/\sqrt{V}}\Big{[}1+O\Big{(}\frac{1}{\sqrt{V}}\Big{)}\Big{]}.$ (2.14)
@@
$G_{V}^{(2)}(p_{V})\leq G_{V}^{(2)}(1+sV^{-1/2})\leq e^{bs}\sum_{n=N}^{\lfloor b\sqrt{V}\rfloor}\frac{1}{n^{3/2}}\lesssim\frac{e^{bs}}{\sqrt{N}}.$ (2.17)
@@ 3.1 Bound on $C(n,n+k)$
We use the asymptotic formula (3.5), and use that $w_{k}=1+O(1/k)$ is bounded. The error term in (3.5) is bounded by a constant since $k$ is at most linear in $n$. The factor $e^{a(x)}$ is also bounded by a constant, by Lemma 3.1(i). We therefore only need to estimate $e^{n\varphi(x)}$. Since $0\leq\frac{k}{n}\leq\frac{t^{2}}{3e}$ and $x=1+\frac{k}{n}$, by (3.2) we have $y(x)\leq\sqrt{3(x-1)}=\sqrt{3k/n}\leq t/\sqrt{e}$, so Lemma 3.1(ii) gives
@@ 3.2 Bound on the surplus generating function
$A(n,z)=\frac{1}{n^{n-2}}\sum_{k=0}^{n}C(n,n+k)z^{k+1},\quad B(n,z)=\frac{1}{n^{n-2}}\sum_{k=\lfloor\frac{1}{2}n\rfloor}^{\infty}C(n,n+k)z^{k+1},$ (3.15)
@@ 3.3 Proof for connected subgraphs
$\Delta_{V}^{(1)}\leq C_{b}V^{-1/4}\sum_{n=3}^{\lfloor b\sqrt{V}\rfloor}\binom{V-1}{n-1}\Big{(}\frac{p}{eV}\Big{)}^{n-1}n^{n-1}\leq C_{b}V^{-1/4}\chi_{V}^{t}(p).$ (3.31)
@@
$\Delta_{V}^{(2)}(1+sV^{-1/2})\lesssim\sum_{n=\lceil b\sqrt{V}\rceil}^{V}\binom{V-1}{n-1}\bigg{(}\frac{1+sV^{-1/2}}{eV}\bigg{)}^{n-1}n^{n-1}\exp\Bigl{\{}\frac{1}{5}\frac{n^{3}}{V^{2}}\Bigr{\}}.$ (3.34)
@@
$\Delta_{V}^{(2)}(1+sV^{-1/2})\lesssim\sum_{n=\lceil b\sqrt{V}\rceil}^{V}e^{-(\frac{1}{2}-\frac{1}{5})n^{2}/V}\frac{1}{\sqrt{n}}e^{|s|n/\sqrt{V}}\leq C_{\lvert s\rvert}b^{-1/2}V^{1/4}.$ (3.35)
@@
$\Delta_{V,0}^{(2)}(1+sV^{-1/2})\lesssim\sum_{n=\lceil b\sqrt{V}\rceil}^{V}e^{-(\frac{1}{2}-\frac{1}{5})n^{2}/V}\frac{1}{n^{3/2}}e^{|s|n/\sqrt{V}}\leq C_{\lvert s\rvert}b^{-3/2}V^{-1/4}.$ (3.38)